what is the electric potential given 0.58m away from 18uC charge?





1
Expert's answer
2022-03-16T18:29:56-0400

"V=\frac{kq}{r}"


"V=\frac{9\times10^9\times18\times10^{-6}}{0. 58^2}=4.82\times10^{-5}V"